(no subject)

arenton@uvahef.phys.virginia.edu
Sat, 19 Jun 1999 01:26:33 -0400

Shift Summary for eve shift, 18/jun/99

On Shift:Elliott Cheu, Mike Arenton, and others,
particularly Hogan.

Hours of beam delivered at all intensities: 5.5
Hours of beam delivered at nominal intensity: 5.5
Hours of E832 data (special runs not included): 0.75
Hours of special runs (laser scans, muon runs...):2.5

Accounting of lost time: access to fix dynode channel 581

========================================================================
Runs ended this shift:

Run 13652-13659 - BEAM_MUON runs for DC plateau

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift:

before beam arrived, several to work on Etot controller boards
and to adust DC offsets in BA.

2hr access to track down problem with block 581 dynode.

========================================================================
Overview of shift:
Shift started with beam off after rad surveys.

Vacuum system PLC had stopped, finally Doug and Bob Sanders were
able to reload the system and restart it (after finding a key needed to
load the software). However one of the diffusion pumps does not come
on. Currently we are running with only one diffusion pump, pressures
are fine with one pump.

J.Hamm replaced controller board in one of Etot crates several
time, finally returning to the original board.

S. Turner adjusted DC offsets in BA, and swapped an ADC used for
RC10 to repair a bad channel.

J.Graham made changes to ONSPOOL.

Power supply group worked on NM2EU,

DC group decides to run DC1 with downstream window at same voltage
as cathodes. A series of runs to plateau DC1 and DC2 were taken after
beam returned at 18:30.

First real data run began at 2100, at 21:45 discoverd that CsI 851
had no HCC bits set. Made an access. First replaced Etot board, that
didn't fix the problem. Then looked at dynode signal into Etot board
on scope, it wasn't there. Hogan went into blockhouse to fix dynode.
Final result was that channel was OK after unplugging and reconnecting
it in the blockhouse.

Out of access at 23:30, ops have trouble getting interlocks reset.
(as this is written 10 minutes into next shift beam has just returned)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13660 16 21:10 burst of ktevana errors, several KFREND's

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Vacuum is running with only one diffusion pump. Probably this
can wait until Monday.

We are starting to fill out run sheets.

NM4AN was reversed and muon runs taken on this shift, no
need to reverse on owl shift.